# Multiscale scattered data analysis in samplet coordinates

### v1

@@ 3. Condition number bound for generalized Vandermonde matrices
We now show that the condition number $\operatorname{cond}_{2}$ of the generalized Vandermonde matrix $\boldsymbol{K}_{\ell,\ell}$ of the scaled kernel can be bounded by a constant, independently of the level $\ell$. A similar result in the context of the multiscale method for compactly supported RBFs has been derived in [18] using techniques of [15, 16]. We will reuse these techniques, with appropriate modifications. Furthermore, for a lighter notation, we will drop the subscripts $\ell$ in this section.
It goes without saying that, such a bound can not be established for general sets of sites $X$. Hence, for this section, we assume that $X$ is quasi-uniform and the parameters of the multilevel method are chosen as in Theorem 2.1.
We remark that the algebraic decay of the Fourier transform of the Matérn kernel (2) immediately yields for the scaled RBF $\Phi_{\delta}$ that
(10) $\widehat{\Phi_{\delta}}(\boldsymbol{\omega})=(1+\delta^{2}\|\boldsymbol{\omega}\|_{2}^{2})^{-\nu-\frac{d}{2}},\quad\boldsymbol{\omega}\in\mathbb{R}^{d}.$
We will bound the smallest and largest eigenvalue of the matrix ${\boldsymbol{K}}$ separately. We follow the ideas of [18, 15, 16].
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain and $X\subseteq\Omega$ be a quasi-uniform set of sites with fill-distance $h_{X,\Omega}$ and separation radius $q_{X}$. Let $\Phi$ be the Matérn kernel and Let $\Phi_{\delta}=\delta^{-d}\Phi(\cdot/\delta)$ be the rescaled kernel where $\delta=\eta h_{X,\Omega}$ with overlap parameter $\eta>1$. Then there exists a constant $C=C(\Phi,d,s)$ such that the bound for the smallest eigenvalue $\lambda_{min}$ of $\boldsymbol{K}=[\Phi_{\delta}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})]_{i,j}$
$\lambda_{min}({\boldsymbol{K}})\geq C(\eta c_{\textnormal{qu}})^{d-2s}\delta^{-d}$
holds. The largest eigenvalue of ${\boldsymbol{K}}$ can be bounded from above by
$\lambda_{max}({\boldsymbol{K}})\leq\widetilde{C}\delta^{-d},$
with a constant $\widetilde{C}=C(\Phi,d,\nu,\eta,c_{\textnormal{qu}})$.
@@
The lower bound follows immediately from [18, Proof of Proposition 3.4].
We only need to prove the upper bound for $\lambda_{max}({\boldsymbol{K}})$. We use Gershgorin’s circle theorem, see, e.g. [23, Theorem 5.1] for $\widetilde{\boldsymbol{K}}$. Since the diagonal elements in this matrix are $\Phi(\boldsymbol{0})$, independent of the line $j$, we can, without loss of generality, set $j=1$ and $\boldsymbol{x}_{1}=\boldsymbol{0}$. We have
@@
see, e.g., [20, Proof of Theorem 12.3]. Inserting this into (11) yields
$|\lambda_{max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{k=2}^{N}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
$\leq\sum_{n=1}^{\infty}\sum_{\boldsymbol{x}_{k}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
(12) $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}.$
Next, we bound $\sup_{\boldsymbol{x}\in E_{n}}|\Phi({\boldsymbol{x}}/{\delta})|$. We use the specific form of $\Phi$, cp. (1), and the fact that the function $r\mapsto r^{\beta}K_{\beta}(r)$ is non-increasing on $(0,\infty)$ for any $\beta\in\mathbb{R}$, see [20, Corollary 5.12]. This yields
$\frac{\Gamma(\nu)}{2^{1-\nu}}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$ $=\sup_{\boldsymbol{x}\in E_{n}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}$
$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}K_{\nu-\frac{d}{2}}\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}$
$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\nu-\frac{d}{2}}\sqrt{2\pi}\bigg{(}\frac{\delta}{nq_{X}}\bigg{)}^{\frac{1}{2}}\exp\bigg{(}-\frac{nq_{X}}{\delta}\bigg{)}\exp\Bigg{(}\bigg{(}\nu-\frac{d}{2}\bigg{)}^{2}\frac{2\delta}{2nq_{X}}\Bigg{)},$
where we used the asymptotic behavior [20, Lemma 5.13] of the modified Bessel function to arrive at the last inequality. Inserting this into (3), using the quasi-uniformity of $X$ and the coupling of $\delta$ to $h_{X,\Omega}$, as in Theorem 2.1, yields
$|\lambda_{max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$
$\leq\sum_{n=1}^{\infty}3^{d}n^{d-1}\sqrt{2\pi}\eta^{\frac{d}{2}-\nu}(\eta c_{\text{qu}})^{\frac{1}{2}}n^{\nu-\frac{d}{2}-\frac{1}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{(\nu-\frac{d}{2})^{2}\eta c_{q}u}{2n}\bigg{)}$
$\leq\sqrt{2\pi}3^{d}\eta^{\frac{d}{2}-\nu}(\eta c_{\text{qu}})^{\frac{1}{2}}\sum_{n=1}^{\infty}n^{\nu-\frac{d}{2}-\frac{3}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{(\nu-\frac{d}{2})^{2}\eta c_{q}u}{2n}\bigg{)}.$
Since the series on the right hand side of the estimate converges, this yields the claimed bound for $\lambda_{max}$. ∎
The derived bounds on $\lambda_{min}$ and $\lambda_{max}$ lead to the boundedness of the condition number of the generalized Vandermonde matrix for every level, uniformly in the number of data sites and the scaling parameter.

### v2

@@ 3. Condition number bound for generalized Vandermonde matrices
We now show that the condition number of the generalized Vandermonde matrices $\boldsymbol{K}_{\ell,\ell}$, $\ell=1,\ldots,L$, of the scaled kernel can be bounded by a constant, independently of the level $\ell$. A similar result in the context of the multiscale method for compactly supported RBFs has been derived in [26] using techniques of [22, 24]. We will reuse these techniques, with appropriate modifications. Furthermore, for a lighter notation, we will drop the subscripts $\ell$ in this section.
It goes without saying that, such a bound can not be established for general sets of data sites $X$. Hence, for this section, we assume that $X$ is quasi-uniform and the parameters of the multilevel method are chosen as in Theorem 2.1.
We remark that the algebraic decay of the Fourier transform of the Matérn kernel (2) immediately yields for the scaled RBF $\Phi_{\delta}$ that $\mathcal{F}{\Phi_{\delta}}(\boldsymbol{\omega})=(1+\delta^{2}\|\boldsymbol{\omega}\|_{2}^{2})^{-\theta},\ \boldsymbol{\omega}\in\mathbb{R}^{d}.$ We will bound the smallest and largest eigenvalue of the matrix ${\boldsymbol{K}}$ separately. We follow the ideas of [26, 22, 24].
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain and $X\subseteq\Omega$ be a quasi-uniform set of data sites with fill-distance $h_{X,\Omega}$ and separation radius $q_{X}$. Let $\Phi$ be the Matérn kernel with smoothness parameter $\theta$. Let $\Phi_{\delta}=\delta^{-d}\Phi(\cdot/\delta)$ be the rescaled kernel where $\delta=\eta h_{X,\Omega}$ with overlap parameter $\eta>1$. Then there exists a constant $C={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}C(\Phi,d)}$ such that the the smallest eigenvalue $\lambda_{\min}$ of $\boldsymbol{K}=[\Phi_{\delta}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})]_{i,j}$ satisfies $\lambda_{\min}({\boldsymbol{K}})\geq C(\eta c_{\textnormal{qu}})^{d-2\theta}\delta^{-d},$ while the largest eigenvalue of ${\boldsymbol{K}}$ can be bounded from above by $\lambda_{\max}({\boldsymbol{K}})\leq\widetilde{C}\delta^{-d},$ with a constant $\widetilde{C}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}C(\Phi,d,\eta,c_{\textnormal{qu}})}$.
@@
The lower bound immediately follows from [26, Proof of Proposition 3.4]. Hence, we only prove the upper bound for $\lambda_{\max}({\boldsymbol{K}})$. To this end, we apply Gershgorin’s circle theorem, see, e.g., [31, Theorem 5.1] to $\widetilde{\boldsymbol{K}}$. Since the diagonal elements of this matrix are $\Phi(\boldsymbol{0})$, independently of the row $i$, we can, without loss of generality, set $i=1$ and $\boldsymbol{x}_{1}=\boldsymbol{0}$. We have
@@
see, e.g., [28, Proof of Theorem 12.3]. Inserting this into (12) yields
(13) $|\lambda_{\max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{k=2}^{N}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}\leq\sum_{n=1}^{\infty}\sum_{\boldsymbol{x}_{k}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}_{k}}{\delta}\bigg{)}\bigg{|}$
$\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}.$
Next, we bound $\sup_{\boldsymbol{x}\in E_{n}}|\Phi({\boldsymbol{x}}/{\delta})|$. We use the specific form of $\Phi$, cp. (1), and the fact that the function $r\mapsto r^{\beta}K_{\beta}(r)$ is non-increasing on $(0,\infty)$ for any $\beta\in\mathbb{R}$, see [28, Corollary 5.12]. This yields
$\frac{\Gamma\left(\theta-\frac{d}{2}\right)}{2^{1-\theta+\frac{d}{2}}}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}=\sup_{\boldsymbol{x}\in E_{n}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}^{\theta-\frac{d}{2}}K_{\theta-\frac{d}{2}}\bigg{(}\frac{\|\boldsymbol{x}\|_{2}}{\delta}\bigg{)}$
$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\theta-\frac{d}{2}}K_{\theta-\frac{d}{2}}\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}$
$\leq\bigg{(}\frac{nq_{X}}{\delta}\bigg{)}^{\theta-\frac{d}{2}}\sqrt{2\pi}\bigg{(}\frac{\delta}{nq_{X}}\bigg{)}^{\frac{1}{2}}\exp\bigg{(}-\frac{nq_{X}}{\delta}\bigg{)}\exp\Bigg{(}\bigg{(}\theta-\frac{d}{2}\bigg{)}^{2}\frac{2\delta}{2nq_{X}}\Bigg{)},$
where we used the asymptotic behavior [28, Lemma 5.13] of the modified Bessel function to arrive at the last inequality. Inserting this into (13), using the quasi-uniformity of $X$ and the coupling of $\delta$ to $h_{X,\Omega}$, as in Theorem 2.1, yields
$|\lambda_{\max}(\widetilde{\boldsymbol{K}})-\Phi(\boldsymbol{0})|$ $\leq\sum_{n=1}^{\infty}\#\{\boldsymbol{x}_{k}\in E_{n}\}\sup_{\boldsymbol{x}\in E_{n}}\bigg{|}\Phi\bigg{(}\frac{\boldsymbol{x}}{\delta}\bigg{)}\bigg{|}$
$\leq\sum_{n=1}^{\infty}3^{d}n^{d-1}\sqrt{2\pi}\eta^{\frac{d}{2}-\theta}(\eta c_{\text{qu}})^{\frac{1}{2}}n^{\theta-\frac{d}{2}-\frac{1}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{\left(\theta-\frac{d}{2}\right)^{2}\eta c_{\text{qu}}}{n}\bigg{)}$
$\leq\sqrt{2\pi}3^{d}\eta^{\frac{d}{2}-\theta}(\eta c_{\text{qu}})^{\frac{1}{2}}\sum_{n=1}^{\infty}n^{\theta-\frac{d}{2}-\frac{3}{2}}\exp\bigg{(}-\frac{n}{\eta}+\frac{\left(\theta-\frac{d}{2}\right)^{2}\eta c_{\text{qu}}}{n}\bigg{)}.$
Since the series on the right hand side of the estimate converges, this yields the claimed bound for $\lambda_{\max}$. ∎
The derived bounds on $\lambda_{\min}$ and $\lambda_{\max}$ lead to the boundedness of the condition number of the generalized Vandermonde matrices ${\boldsymbol{K}}_{\ell,\ell}$ for $\ell=1,\ldots,L$, uniformly in the number of data sites and the scaling parameter. In addition, by a suitable diagonal scaling, also the lower triangular system in (10) becomes well conditioned. We have the following